OpenStudy (anonymous):

Find the center of a circle with equation x^2+y^2-6x+4y-4=0

OpenStudy (jlastino):

Use the "completing the square" method group similar terms and transpose all constants to the right side of the equation first...

OpenStudy (anonymous):

what are constants

OpenStudy (anonymous):

A.) (3,2) B.) (3,-2) C.) (-3,2) D.) (-3,-2)

OpenStudy (jlastino):

Those without variables( 2, 3, 4 )...are you in a hurry? do you need the answer right away?

OpenStudy (anonymous):

im taking my graduation test but i have time

OpenStudy (anonymous):

im not very good at learning something over the computer

OpenStudy (jlastino):

Oh, i see..so ,what is the constant in the equation? x^2+y^2-6x+4y-4=0

OpenStudy (anonymous):

um 6x and 4y ?? can there only be one?

OpenStudy (anonymous):

or was the constant the number that didnt have a letter in it?? i cant remember

OpenStudy (jlastino):

no....constant are only composed of numbers (they don't have x's or y's) well at the state of our equation there is only one but there will be later when we expand it

OpenStudy (jlastino):

i mean yeah, those without letters are constants

OpenStudy (anonymous):

ok so its -4

OpenStudy (jlastino):

yes , now when you transpose it to the other side, its sign changes

OpenStudy (anonymous):

ok so its a positive 4 now. do i just move the constants to the right side?

OpenStudy (anonymous):

cause its already on the right side

OpenStudy (jlastino):

yup can you show me the equation?

OpenStudy (anonymous):

um.. X^2+y^2-6x+4y+4=0

OpenStudy (jlastino):

with the constants on the right side..

OpenStudy (anonymous):

it is on the right side

OpenStudy (jlastino):

oh what i meant was at the right side of the = sign (that's the right side of the equation)

OpenStudy (anonymous):

so it would be =0+4 on the right side? im sorry im really bad at this

OpenStudy (jlastino):

yeah but you don't need to write the 0 since anything added to 0 is equal to itself...now what is our equation so far?

OpenStudy (anonymous):

oh ok x^2+y^2-6x+4y=4

OpenStudy (jlastino):

correct :D the next thing to do is group the x and y's together inside parentheses ( the one's with x go here ) + ( the ones with y here ) =4

OpenStudy (anonymous):

(x^2-6x) + (y^2+4y)=4

OpenStudy (jlastino):

ok we're almost done . we need to complete the square of the terms inside the parentheses, to do that that divide the second term by two and square it, then add it inside the parentheses for each. also add that to the right side example (x^2-2x+ 1) - 1 because -2/2 =-1 and -1^2 =1 (y^2-8x+16) 16 because -8/2 =-4 and 4^2= 16 (x^2-2x+1) +(y^2-8x+16)= 3 +1+16

OpenStudy (anonymous):

3+2=4 ?

OpenStudy (anonymous):

So, (3,2) ?

OpenStudy (jlastino):

Well its close but here's how you do it (x^2-6x+9) +(y^2+4y+4) = 4+9+4 ^ (x-3)^2 + (y+2)^2= 17 So the answer is 3,-2

OpenStudy (anonymous):

oh ok thanks :)